# VUV Reflectance Measurements for Materials Relevant to Argon and Xenon Experiments

### v1

@@ Abstract
Accurate knowledge of material reflectance in the vacuum ultraviolet (VUV) range is crucial for optimizing photon detection in noble gas detectors such as DUNE. Despite its importance, reflectance values for detector materials in the VUV region remain poorly characterized, with literature values showing significant variation depending on surface termination and finish. We present an angular-resolved reflectance measurement system developed at IFIC that operates in a gaseous argon atmosphere, enabling realistic measurements of detector materials under controlled conditions. The setup couples a deuterium lamp to a monochromator and employs a motorized PMT rotating around the sample to measure reflected light distributions across a wide angular range. We have characterized two key DUNE materials—aluminum field cage profiles and stainless steel cryostat membranes—in both the UV-VIS (300–500 nm) and VUV (128–200 nm) ranges. In the UV-VIS region, we confirm literature values of approximately 60% reflectance for aluminum and 40% for stainless steel. Preliminary VUV measurements at 45° angle of incidence yield reflectance values of 10–15% for both materials, significantly lower than their UV-VIS counterparts. The reflected light distributions exhibit a mixed character between specular and diffuse reflection. These results have direct implications for detector simulations and light yield predictions in next-generation experiments.
@@ 1 Introduction
Scintillation light in noble element detectors serves multiple critical functions: it provides a trigger signal for the time projection chambers as in DUNE [3], enables precise timing reconstruction ($t_{0}$), and allows for calorimetric reconstruction of deposited energy. Reliable calorimetric reconstruction depends critically on accurate knowledge of the light yield through the active volume, defined as the number of photons detected per unit deposited energy (typically expressed in photons/MeV). To achieve this accuracy, a comprehensive understanding of photon propagation within the detector is essential.
A key aspect often overlooked in detector design and simulation is the role of material reflectance. The surfaces of the main detector components can reflect a significant fraction of scintillation photons back into the active volume. Poor characterization of reflectance properties can introduce a significant bias in the light yield determination. In particular, we demonstrated in [5] that considering a reflectance of 40% for the liquid-argon scintillation light on the aluminium field cage and stainless steel cryostat walls of ProtoDUNE Dual-Phase [1] would lead to a bias of 40% in the determination of the simulated light yield with respect to the null reflectance case. Despite this sensitivity, reflectance data for detector materials in the vacuum ultraviolet (VUV) range—where argon and xenon scintillation peaks occur—remain poorly characterized in the literature, with published values showing substantial variation depending on surface termination and finish.
@@ 2 Experimental setup
Our measurement system couples a deuterium lamp (model McPherson 621) and a tungsten lamp (model McPherson 634) to a McPherson 302/234 monochromator. This device incorporates an aberration-corrected concave holographic diffraction grating of 1200 grooves/mm, in a Seya-Namioka optical design. The grating is equipped with a wavelength-optimized reflective coating of aluminium, enhanced with magnesium fluoride for VUV throughput. This enables a wavelength selection range of 115-550 nm. The monochromator output is coupled to a collimator and a methacrylate black box where the measurement is performed. The entire system is maintained in a pure GAr atmosphere within a dedicated dark room to allow for measurements in the VUV region. This atmosphere choice unlike traditional vacuum measurements serves two key purposes: (i) it provides the optical transmission necessary for VUV measurements while enabling flexibility to fully instrument the black box and (ii) it protects the monochromator grating from contamination and outgassing.
The angular-resolved measurement capability is provided by a 3D-printed assembly comprising three automated motorized stages. Motor 1 allows sample exchange, Motor 2 rotates a VUV-sensitive PMT (Hamamatsu R6836-Y004) around the sample to set the angle of reflection, and Motor 3 rotates the sample holder to set the angle of incidence (AOI). A rectangular mask (2x20 mm2 opening providing 1.6° angular resolution) placed in front of the PMT detector enhances angular resolution and defines the collection solid angle. A picture and a diagram of the setup are shown in Figure 1.
We measured two primary materials from the DUNE prototypes: (i) aluminium field cage profiles and (ii) stainless steel (SS) cryostat membranes. Calibration samples—including a UV mirror and a UV beam splitter with known reference reflectance values—were employed to validate the measurement methodology across the UV-VIS range.
@@ 3 Methodology
For each wavelength and at a fixed angle of incidence, we record the angular distribution of reflected light by rotating the PMT across a defined angular range (typically $\pm 40^{\circ}$ from the specular direction). Normalization measurements are acquired at each wavelength both before and after sample measurements to account for variations in the deuterium lamp intensity and GAr atmosphere transmission stability (typically maintained to less than 6% in all measurements).
The reflected light distributions exhibit a mixed character between specular and diffuse reflection, well-described by the following equation, inspired by the shading-function model described in [4]:
$I(\theta)=\sum_{i=1}^{2}A_{i}cos^{n_{i}}(\theta-x_{0})$
where $n$ parameterizes surface roughness (higher $n$ indicates more specular behavior), $\theta_{0}$ is the specular direction, $A$ is the normalization amplitude, and $i$ is the summation index. This functional form successfully captures the asymmetric beam profiles observed in our data. An example fit to data with the resulting parameter values is shown in the left panel of Figure 2.
Figure 2: Left: Angular-resolved reflectance intensity for the stainless-steel sample at an angle of incidence of 45º and a wavelength of 300 nm. Right: Reflectance of the $S$ and Al samples.
The integrated (or hemispherical) reflectance is computed by summing the angular-resolved measurements and normalizing to the lamp spectrum. For metallic samples (Al and SS), we apply a correction assuming radial symmetry to account for the light scattered at wide angles beyond the PMT acceptance. This "ring model" divides the emitted light pattern into concentric rings and integrates their contribution, accounting for the mask geometry, to estimate the total emitted power.
@@ 4 Results
4.1 UV-VIS Measurements (300–500 nm)
The right panel of Figure 2 shows the reflectance measurements in the UV-VIS range. They show excellent agreement with the expected values for the reference samples: the UV mirror (Thorlabs PF1011-F01) and beamsplitter (Thorlabs BSW20) reproduce their manufacturer specifications. For DUNE materials, aluminum exhibits approximately 60% reflectance across the UV-VIS range, while stainless steel shows approximately 40%, consistent with literature values, but in this case, for real detector samples.
@@ 4.2 VUV Measurements (128–200 nm) in GAr
Despite the reduced photon flux and lower detector sensitivity in the VUV, we successfully detected reflected light from both materials, as shown in Figure 3. The angular distributions at VUV wavelengths remain qualitatively similar across 128, 150, 175, and 200 nm, with stainless steel exhibiting more specular behaviour compared to aluminium.
Figure 3: Angular-resolved reflected intensity of the $S$ and Al samples at VUV wavelengths.
@@ 4.3 Integrated reflectance
At an angle of incidence of 45 in the VUV range, preliminary reflectance values are 10–15% for both aluminium and stainless steel, significantly lower than their UV-VIS counterparts, see Figure 4 . This substantial drop at shorter wavelengths reflects the wavelength-dependent optical properties of these materials and has important consequences for detector design.
Figure 4: Reflectance of the $S$ and Al samples in the VUV and UV/Visisble range.
@@ 5 Discussion
The measured angular distributions reveal a mixed specular-diffuse character for both aluminium and stainless steel surfaces. This finding has important implications for photon transport modelling: simple assumptions of purely specular or purely Lambertian reflection inadequately describe the true optical behaviour. Implementation of realistic angular distributions derived from the shading-function model [4] parametrization into Monte Carlo simulations should improve predictions of light yield and photon detection efficiency.
The integrated reflectance values measured in the VUV range (10–15% at 45° AOI) are lower than both our UV–VIS measurements and the reflectance values typically assumed in DUNE-VD simulations, where about 70% is considered for aluminium and 30-40% for stainless steel [2]. This discrepancy implies that the contribution of field-cage and cryostat surfaces to the overall light yield should be carefully re-evaluated.
@@ 6 Conclusions and future steps
We have established a dedicated facility at IFIC for angular-resolved reflectance measurements in the VUV range using a GAr atmosphere, enabling realistic measurements of detector materials under controlled conditions. Preliminary results indicate reflectance values in the range of 10–15% for aluminum and stainless steel at an angle of incidence of 45 in the VUV, with mixed angular distributions best described by a shading-function model [4]. These measurements provide a foundation for improved detector simulations and will guide optimization of light collection in next-generation noble gas experiments.
@@ Acknowledgments
The authors acknowledge support from Agencia Estatal de Investigación (projects PID2023-147949NB-C52 and JDC2023-051298-I) and G. Valenciana (ASFAE/2022/029 and ASFAE/2022/028).

### v2

@@ Abstract
Accurate knowledge of material reflectance in the vacuum ultraviolet (VUV) range is crucial for optimizing photon detection in noble gas detectors such as DUNE. Despite its importance, reflectance values for detector materials in the VUV region remain poorly characterized, with literature values showing significant variation depending on surface termination and finish. An angular-resolved reflectance measurement system developed at the Instituto de Física Corpuscular at Valencia (IFIC) that operates in a gaseous argon atmosphere is presented, enabling realistic measurements of detector materials under controlled conditions. The setup couples a deuterium lamp to a monochromator and employs a motorized PMT rotating around the sample to measure reflected light distributions across a wide angular range. We have characterized two key DUNE materials—aluminum field cage profiles and stainless steel cryostat membranes—in both the UV-VIS (300–500 nm) and VUV (128–200 nm) ranges. In the UV-VIS region, we confirm literature values of approximately 60% reflectance for aluminum and 40% for stainless steel. Preliminary VUV measurements at 45° angle of incidence yield reflectance values of 10–15% for both materials, significantly lower than their UV-VIS counterparts. The reflected light distributions exhibit a mixed character between specular and diffuse reflection. These results have direct implications for detector simulations and light yield predictions in next-generation experiments.
@@ 1 Introduction
Scintillation light in noble element detectors serves multiple critical functions: it provides a trigger signal for the liquid-argon time projection chambers as in the Deep Underground Neutrino Experiment (DUNE) [3], enables precise timing reconstruction ($t_{0}$), and allows for calorimetric reconstruction of deposited energy. Reliable calorimetric reconstruction depends critically on accurate knowledge of the light yield through the active volume, defined as the number of photons detected per unit deposited energy (typically expressed in photons/MeV). To achieve this accuracy, a comprehensive understanding of photon propagation within the detector is essential.
A key aspect often overlooked in detector design and simulation is the role of material reflectance. The surfaces of the main detector components can reflect a significant fraction of scintillation photons back into the active volume. Poor characterization of reflectance properties can introduce a significant bias in the light yield determination [13]. In particular, it has been demonstrated in [15] that considering a reflectance of 40% for the liquid-argon scintillation light on the aluminium field cage and stainless steel cryostat walls of ProtoDUNE Dual-Phase [1] would lead to a bias of 40% in the determination of the simulated light yield with respect to the null reflectance case. Despite this sensitivity, reflectance data for detector materials in the vacuum ultraviolet (VUV) range—where argon and xenon scintillation peaks occur—remain poorly characterized in the literature, with published values showing substantial variation depending on surface termination and finish.
@@ 2 Experimental setup
The measurement system couples a deuterium lamp (model McPherson 634 [11]) and a tungsten lamp (model McPherson 621 [10]) to a McPherson 302/234 monochromator [9]. This device incorporates an aberration-corrected concave holographic diffraction grating of 1200 grooves/mm, in a Seya-Namioka optical design. The grating is equipped with a wavelength-optimized reflective coating of aluminium, enhanced with magnesium fluoride. This enables a wavelength selection range of 115-550 nm. The monochromator output is coupled to a collimator and a methacrylate black box where the measurement is performed. When measuring in the VUV range, the entire system is maintained in a pure GAr atmosphere (Air Liquid ALPHAGAZ [7] of 99.999% purity) within a dedicated dark room to allow for measurements in the VUV region. The black box is continuously purged with pure GAr and kept at a slight overpressure of about 100 mbar, monitored by a pressure sensor. This atmosphere choice unlike traditional vacuum measurements serves two key purposes: (i) it provides the optical transmission necessary for VUV measurements while enabling flexibility to fully instrument the black box and (ii) it protects the monochromator grating from contamination, as outgassing in vacuum-based VUV systems can lead to the deposition of carbonaceous films on optical elements and degrade their performance over time [6, 5].
The angular-resolved measurement capability is provided by a 3D-printed assembly comprising three automated motorized stages. Motor 1 allows sample exchange, Motor 2 rotates a VUV-sensitive PMT (Hamamatsu R6836-Y004) around the sample to set the angle of reflection, and Motor 3 rotates the sample holder to set the angle of incidence (AOI). A rectangular mask ($2\times 20$ mm2 opening providing 1.6° angular resolution) placed in front of the PMT detector enhances angular resolution and defines the collection solid angle. A picture and a diagram of the setup are shown in Fig. 1. Since the monochromator provides a high light intensity, the PMT is operated without gain by measuring the first-dynode current with a picoammeter. A 100 V bias between the photocathode and the first dynode ensures full collection efficiency.
Two materials from the DUNE prototypes are measured: (i) 6101 aluminium sample (Al) extracted from a field cage profile and (ii) 304L stainless steel sample (SS) from the cryostat inner membrane [4]. Calibration samples—including a UV mirror and a UV beam splitter with known reference reflectance values—were employed to validate the measurement methodology across the UV-VIS range.
@@ 3 Methodology
For each wavelength and at a fixed angle of incidence, the angular distribution of reflected light is recorded by rotating the PMT across a defined angular range (typically $\pm 40^{\circ}$ from the specular direction. Direct-light normalization measurements (when the PMT is around $90^{\circ}$ in Fig. 1) are acquired at each wavelength both before and after sample measurements to account for variations in the deuterium lamp intensity and GAr atmosphere transmission stability (typically maintained to less than 6% in all measurements). The uncertainty introduced by this variation is added in quadrature for each measurement.
The reflected light distributions exhibit a mixed character between specular and diffuse reflection, well described by the following equation, inspired by the shading-function model described in [14]: $I(\theta)=\sum_{i=1}^{2}A_{i}cos^{n_{i}}(\theta-x_{0})$, where $n$ parameterizes surface roughness (higher $n$ indicates more specular behavior), $x_{0}$ is the specular direction, $A$ is the normalization amplitude, and $i$ is the summation index. This functional form successfully captures the beam profiles observed in our data. An example fit to data with the resulting parameter values is shown in the left panel of Fig. 2. In general SS shows a more specular behavior than Al ($n$ parameter: ~1000 vs. ~55 at 300-500 nm).
Figure 2: Left: Angular-resolved reflectance intensity for the stainless-steel sample at an angle of incidence of $45^{\circ}$ and a wavelength of 300 nm in air. Right: Reflectance of the $S$ and Al samples in the UV/Visible range in air.
The integrated (or hemispherical) reflectance is computed by summing the angular-resolved measurements (when the PMT is around $0^{\circ}$ in Fig. 1) and normalizing to direct light measurement (when the PMT is around $90^{\circ}$ in Fig. 1): $R=\sum_{\theta=-40^{\circ}}^{40^{\circ}}I_{\theta}/\sum_{\theta=80^{\circ}}^{100^{\circ}}I_{\theta}$. For metallic samples (Al and SS), which exhibit more diffuse reflection, the azimuthal scan covers the full range of interest ($\pm 40^{\circ}$) with PMT rotation, while vertical acceptance is limited by the mask height (2 cm or $\pm$8). Radial symmetry is assumed to estimate the uncollected light in the vertical direction.
@@ 4 Results
4.1 UV-Visible Measurements (300–500 nm) in air
The right panel of Fig. 2 shows the reflectance measurements in the UV-VIS range in air. They show excellent agreement with the expected values for the reference samples: the UV mirror (Thorlabs PF1011-F01 [17]) and beamsplitter (Thorlabs BSW20 [16]) reproduce their manufacturer specifications. For DUNE materials, aluminum exhibits approximately 60% reflectance across the UV-VIS range, while stainless steel shows approximately 40%, consistent with literature values, but in this case, for real detector samples [13, 12, 8, 18].
@@ 4.2 VUV Measurements (128–200 nm) in GAr
Despite the reduced photon flux and lower detector sensitivity in the VUV, reflected light is successfully detected from both materials, as shown in Fig. 3. The angular distributions at VUV wavelengths (128–200 nm) remain qualitatively similar, with SS sample showing more specular behavior than Al sample considered in these tests (higher $n$ parameter: ~600 vs. ~80 at 128 nm), and with small wavelength dependence.
Figure 3: Left: Angular-resolved reflected intensity of the $S$ and Al samples at VUV wavelengths. Right: Reflectance of the $S$ and Al samples in the VUV and UV/Visible range.
@@ 4.3 Integrated reflectance
At an angle of incidence of 45 in the VUV range, preliminary reflectance values are 10% (15%) for SS (Al) samples, significantly lower than their UV-VIS counterparts, as it is shown in the right panel of Fig. 3. Al sample is slightly more reflective than SS sample, but far below some literature values [18], highlighting the strong dependence on surface finish and the need for sample-specific measurements. This substantial drop at shorter wavelengths reflects the wavelength-dependent optical properties of these materials and has important consequences for detector design.
@@ 5 Discussion
The measured angular distributions reveal a mixed specular-diffuse character for both aluminium and stainless steel surfaces. This finding has important implications for photon transport modelling: simple assumptions of purely specular or purely Lambertian reflection inadequately describe the true optical behaviour. Implementation of realistic angular distributions derived from the shading-function model [14] parametrization into Monte Carlo simulations should improve predictions of light yield.
The integrated reflectance values measured in the VUV range are lower than the reflectance values typically assumed in DUNE simulations, where about 70% is considered for aluminium and 30-40% for stainless steel [2]. This discrepancy implies that the contribution of field-cage and cryostat surfaces to the overall light yield should be carefully re-evaluated.
@@ 6 Conclusions and future steps
A dedicated facility at IFIC for angular-resolved reflectance measurements in the VUV range using a GAr atmosphere as been stablished, enabling realistic measurements of detector materials under controlled conditions. Preliminary results indicate reflectance values in the range of 10–15% for aluminum and stainless steel at an angle of incidence of 45 in the VUV, with mixed angular distributions best described by a shading-function model [14]. These measurements provide a foundation for improved detector simulations and will guide optimization of light collection in next-generation noble gas experiments.
@@ Acknowledgments
The authors acknowledge support from Agencia Estatal de Investigación (projects PID2023-147949NB-C52 and JDC2023-051298-I) and Generalitat Valenciana (ASFAE/2022/029 and ASFAE/2022/028).